OpenStudy (anonymous):

Mike observed that 75% of the students of a school liked skating. If 35 students of the school did not like skating, the number of students who liked skating was ______. (only put numeric values, no other symbols)

OpenStudy (danjs):

35 students not like skating, that is (100-75)% = 25% of the total students

OpenStudy (anonymous):

so my answer is (100-75)% =25

OpenStudy (danjs):

35 is 25% of total 35 = 0.25*(total)

OpenStudy (danjs):

The total # of students is 140

OpenStudy (danjs):

35 Not like skating 140 - 35 = 105 students like skating

OpenStudy (anonymous):

so 105 students like skating

OpenStudy (danjs):

yes

OpenStudy (anonymous):

okay thx and i was trying to see what u were doing but i understand thx for your help